A chemist wants to make 52 mL of a 17% acid solution by mixing a 11% acid solution and an 19% acid solution. How many milliliters of each solution should the chemist use?
Answer by mananth(16946)

percent ---------------- quantity
Acid solution I 19.00% ---------------- x ml
Acid solution II 11.00% ------ 52 - x ml
Mixture 17.00% ---------------- 52
Total 52 ml
19.00% x + 11.00% ( 52 - x ) = 17.00% * 52
19 x + 11 ( 52 - x ) = 884
19 x + 572 - 11 x = 884
19 x - 11 x = 884 - -572
8 x = 312
/ 8
x = 39 ml 19.00% Acid solution I
13 ml 11.00% Acid solution II

m.ananth@hotmail.ca